OpenStudy (anonymous):

y=(x-2/3)(x+4). Find x when y=0

OpenStudy (anonymous):

x= 2/3, -4

OpenStudy (cgreenwade2000):

If y = 0, then x-2/3 = 0 and x+4=0

OpenStudy (anonymous):

how did you get it

OpenStudy (anonymous):

oh so dont multiply the two equations

OpenStudy (cgreenwade2000):

Well if (a+1)(a+2)= 0 Then you can think that if one of them equals 0 then it doesnt matter what the other is. (0 times anything = 0)

OpenStudy (cgreenwade2000):

No you dont have to multiply. so you break that down and say, a+1 =0 because if it does equal zero, a+2 can be anything and the answer will still come to 0.

OpenStudy (anonymous):

oh okay! I have another question kinda similar that I am going to post